Can you please help me solve using the elimination method?
a + 3b = 26
2a - 3b = -2

Answer by 303795(602)

Equation 1 : a + 3b = 26
Equation 2 : 2a - 3b = -2
Add the two equations
3a = 24
a = 8
Substitute a = 8 in one of the equations
8 + 3b = 26
8 + 3b - 8 = 26 - 8
3b = 18
b=6
ie a = 8 and b = 6